# SLOPE and Designing Robust Studies for Generalization

### v1

@@ 3.2 Example: SLOPE for the Mean and the Robustness of a Homogeneous Source Population
Then SLOPE for the mean is the derivative of $\psi^{{\textnormal{mean}}}(Q_{O,X}^{\gamma})$ with respect to $\gamma$, evaluated at $\gamma=0$. Theorem 1 formally states the result under regularity conditions to make this derivative possible.
Theorem 1 (SLOPE for Mean).
@@
Suppose Condition 1 holds with $s(O,\psi^{{\textnormal{mean}}})=O-\psi^{{\textnormal{mean}}}$. Then SLOPE for the mean is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})=\textnormal{E}_{Q_{X}}\{\sigma^{2}(X)\},$ (6)
where $\sigma^{2}(X)=\textnormal{Var}_{P_{O\mid X}}(O\mid X)$ is the conditional variance in the source population.
In words, the SLOPE of the mean is an average of $\sigma^{2}(X)$, the conditional variance of $O$ under $P_{O\mid X}$, over the target population (i.e., $Q_{X}$). When $\sigma^{2}(X)$ is constant across $X$ (i.e., homoskedastic), the mean’s SLOPE only depends on the marginal variance of $O$ in the source population. In other words, under homoskedasticity, SLOPE suggests that a source population with a small variance of $O$ is less sensitive to local violation of exchangeability.
This theoretical insight agrees with existing empirical recommendations for generalizing treatment effects, i.e., under the potential outcome framework in Example 4 with O=Y(1)Y(0)O=Y(1)-Y(0) representing an individual treatment effect. As noted from Tipton and Olsen, (2018), a more homogeneous individual treatment effect (say $O$ is constant in $P$) is generally more robust for generalization than a heterogeneous treatment effect, given all else is equal (i.e., sufficient overlap and covariate balance between the two populations). In Corollary 1, we present one simple mathematical formalization of the empirical insights from Tipton and Olsen, (2018) by using our proposed SLOPE.
Corollary 1 (SLOPE for Robust Source Distributions).
@@
Fix the target distribution $Q_{X}$ and consider two source populations $P$ and $\widetilde{P}$ that satisfy $\sigma^{2}(X)\leq\widetilde{\sigma}^{2}(X)$ almost surely $Q_{X}$, where $\sigma^{2}(X)=\textnormal{Var}_{P_{O\mid X}}(O\mid X)$ and $\widetilde{\sigma}^{2}(X)=\textnormal{Var}_{\widetilde{P}_{O\mid X}}(O\mid X)$. Then
@@
where the equality holds if and only if $\sigma^{2}(X)=\widetilde{\sigma}^{2}(X)$ almost surely $Q_{X}$.
To better illustrate Corollary 1, consider a toy example where $X$ is binary with values $x_{1}$ and $x_{2}$. Condition on $X$, suppose the source is normal with $P_{O\mid X=x_{1}}\sim N(\mu_{1},\sigma_{1}^{2})=N(0,0.5^{2})$ and $P_{O\mid X=x_{2}}\sim N(\mu_{2},\sigma_{2}^{2})$. Fixing $\mu_{2}=0.3$, top panels in Figure 2 show that as $\sigma_{2}^{2}$ increases (i.e., the conditional variance increases), the SLOPE for the mean increases (i.e., the sensitivity increases).
Figure 2: SLOPE of the mean in a toy example with two source distributions, $P_{O\mid X=x_{1}}\sim N(0,0.5^{2})$ and $P_{O\mid X=x_{2}}\sim N(\mu_{2},\sigma_{2}^{2})$. The two left panels fix $\mu_{2}=0.3$ and vary $\sigma_{2}$ and the right panels fix $\sigma_{2}=0.6$ and vary $\mu_{2}$, each with two target populations. Each panel plots the SLOPE in the y-axis against the data generating parameter ($\sigma_{2}$ or $\mu_{2}$) in the x-axis, for SLOPEs of the mean (circles on solid lines) and the median (crossmarks on dotted lines).
Similarly to Corollary 1, for a fixed source population, Corollary 2 states that one target population is more robust than another if the former has a lower measure on covariate values that correspond to a high conditional variance. More concretely, consider again the toy example with $\sigma_{2}=0.6>0.5=\sigma_{1}$. Then $x_{1}$ and $x_{2}$ can be viewed as two subgroups where units with $X=x_{2}$ are more heterogeneous. Consider two target populations $Q$ and $\widetilde{Q}$ where $Q_{X}(x_{2})=0.5$ and $\widetilde{Q}_{X}(x_{2})=0.8$. Since $_{X}(x_{2})<\widetilde{Q}_{X}(x_{2}), QQ$ has a lower proportion of units from the more heterogeneous subgroup than $\widetilde{Q}$, and thus the SLOPE under $Q$ is lower than the SLOPE under $\widetilde{Q}$. As illustrated in Figure 2, when $\sigma_{2}>\sigma_{1}$, the top left panel (i.e., with $Q$) has a smaller SLOPE than the top right panel (i.e., with $\widetilde{Q}$), and vice versa.
Corollary 2 (SLOPE for Robust Target Distributions).
Fix the source distribution $P_{O,X}$ and consider two target distributions $Q_{X}$ and $\widetilde{Q}_{X}$ over a common support $\mathcal{S}_{X}$ such that there exists a subset $\mathcal{S}_{X,1}\in\mathcal{S}_{X}$ that satisfies $Q_{X}(\mathcal{S}_{X,1})\leq\widetilde{Q}_{X}(\mathcal{S}_{X,1})$. If there exists a constant $c$ such that $\sigma^{2}(X)<c$ for $X\in\mathcal{S}_{X,1}$ and $\sigma^{2}(x)>c$ for $x\not\in\mathcal{S}_{X,1}$, then
@@
where the equality holds if and only if $Q_{X}(\mathcal{S}_{X,1})=\widetilde{Q}_{X}(\mathcal{S}_{X,1})$.
@@ 3.3 Example: SLOPE for Median
Similar to the SLOPE of the mean, the SLOPE of the median also depends on the dispersion of the underlying distribution $Q_{O,X}^{0}$. We provide its general form in Theorem 2 along with two special cases.
Theorem 2 (SLOPE for Median).
Suppose Condition 1 holds with $s(O,X,\psi^{{\textnormal{med}}})=\mathds{1}(O\leq\psi^{{\textnormal{med}}})-\mathds{1}(O>\psi^{{\textnormal{med}}})$. Then SLOPE for the median is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})=$ $\dfrac{\textnormal{E}_{Q_{X}}\left[F_{P_{O\mid X}}(m_{1/2}\mid X)\mu(X)\right]-\textnormal{E}_{Q_{O,X}^{0}}\left[O\mathds{1}(O\leq m_{1/2})\right]}{f_{Q_{O}^{0}}(m_{1/2})}.$ (7)
(i) If $P_{O\mid X}$ is symmetric and $\mu(X)=m_{1/2}$ almost surely $Q_{X}$, then (7) simplifies to
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{med}})=\dfrac{m_{1/2}-\textnormal{E}_{Q_{O}^{0}}\left[O\mid O\leq m_{1/2}\right]}{2f_{Q_{O}^{0}}(m_{1/2})}.$ (8)
@@
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})=\textnormal{E}_{Q_{X}}\left[\sigma^{2}(X)\cdot\dfrac{f_{P_{O\mid X}}(m_{1/2})}{\textnormal{E}_{Q_{X}}\left\{f_{P_{O\mid X}}(m_{1/2})\right\}}\right].$
Compared with the mean, the SLOPE for the median crucially depends on the marginal median $m_{1/2}$ and the tail behavior. Therefore, it reflects the spread of $Q_{O,X}^{0}$ in ways that are more complicated than $\sigma^{2}(X)$. For example, in part (i) which induces symmetry, the SLOPE of the median in (8) is the product of two quantities: (a) the difference between the mean and the truncated mean that is lower than the median divided by two, $\left\{m_{1/2}-\textnormal{E}_{Q_{O}^{0}}\left[O\mid O\leq m_{1/2}\right]\right\}/2$, which measures the spread of $Q_{O}^{0}$ and (b) the inverse of the marginal density at the median, $1/f_{P_{O\mid X}}(m_{1/2})$, which is called Tukey’s sparsity (Tukey,, 1965) and measures the inverse of the concentration of $Q_{O}^{0}$. In part (ii) when $P_{O\mid X}$ is Gaussian, the SLOPE of the median becomes a weighted average of the conditional variance $\sigma^{2}(X)$ where the weight is determined by the conditional density $f_{O\mid X}$ evaluated at the median $m_{1/2}$. This form of the median’s SLOPE resembles the mean’s SLOPE (6), which is an (unweighted) average of $\sigma^{2}(X)$.
As the top row of Figure 2 shows, SLOPE for the median also increases when (a) the variance $\sigma_{2}^{2}$ increases, or (b) when the target population assigns a higher proportion for the subgroup with a higher variance. In general for the median, the choice of a robust distribution is more complex since $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{med}})$ depends on $Q_{O,X}^{0}$ not only through the conditional variance $\sigma^{2}(X)$, but also the shape and tail behavior of the conditional density $f_{P_{O\mid X}}$.
@@ 4.1 Relationship Between SLOPE and IF
Another measure of robustness that precedes and complements our work is Hampel, (1974)’s celebrated influence function (IF). In this section, we show how SLOPE is related to IF.
Briefly, the IF of $\psi(\cdot)$ under $Q_{O,X}^{0}$ is a derivative-based local measure of robustness that quantifies the effect of an infinitesimal contamination at the point $(o,x)$ on the estimand $\psi(Q_{O,X}^{0})$,
@@
The term $\delta_{o,x}$ represents the dirac delta function at $(o,x)$. The IF was originally proposed by Hampel, (1974) as the “influence curve” and later called “influence function” in view of its generalization to higher dimensions (Hampel et al.,, 2011).
Both SLOPE and IFs are local measures of robustness. The main difference is that, IFs measure the local change of the target estimand due to contamination of a single data point whereas SLOPE measures the local change of the target estimand due to violation of conditional exchangeability from all data points. Nevertheless, since both can be viewed as evaluations of directional derivatives, SLOPE is equal to a “tilting” of the IF onto a particular direction as the following Theorem shows.
@@
Let $\mu(X)=\textnormal{E}_{P_{O\mid X}}(O\mid X)$ be the conditional mean of $O$ given $X$ on the source population. Then SLOPE can be written as
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)$ $=\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{O\mid X}}\left[\textnormal{IF}(O,X,\psi(Q_{O,X}^{0}))\left\{O-\mu(X)\right\}\mid X\right]\right),$ (10)
whenever either (i) $\psi$ is a Z-functional defined in (2) and Condition 1 hold, or (ii) Conditions 2-3 in the Supplement hold.
Similar to Condition 1, condition (ii) in Theorem 3 ensures the chain rule under Hadamard differentiability holds. In words, equation (10) states that SLOPE is the conditional covariance of the IF and $O-\mu(X)$. As shown in Remark 7 of the Supplement, the term $O-\mu(X)$ is a result of the sensitivity model (3) and can be interpreted as the “direction” of violating conditional exchangeability. From this perspective, SLOPE is the projection of the IF onto the direction in which conditional exchangeability is violated. If the IF is nearly orthogonal to the subspace based on violating conditional exchangeability (or more specifically, $O-\mu(X)$), SLOPE will be close to zero.
In addition to the geometric interpretation of SLOPE, Theorem 3 provides a general formula to derive the SLOPE given an IF. Sections 3.2 and 3.3 derive SLOPEs for the mean and the median using their respective IFs. Also, Section C of the Supplement derives SLOPEs for scale parameters, ordinary least squares (OLS) coefficients, Pearson correlation, L-functionals and some other Z-functionals.
@@ 4.2 Estimation
We briefly discuss two estimators of SLOPE, the weighting estimator and the regression estimator. In short, estimation of SLOPE follows from the theory of M-estimation (e.g., Chapter 5 of Van der Vaart, (2000)) and to avoid overloaded notations, we focus on estimating the SLOPE of Z-functionals defined in equation (2), and defer the implementation details, theoretical results on the asymptotic properties, and simulations to Sections E and G of the Supplement.
Suppose we have $n_{q}$ independent and identically distributed (i.i.d.) samples $X_{i}\sim Q_{X}$ with $i=1,\cdots n_{q}$ and another independent, $n_{p}$ i.i.d. samples $(X_{i},O_{i})\sim P_{O,X}$ with $i=n_{q}+1,\cdots,n_{p}+n_{q}=n$. Under part (i) of Condition 1, the SLOPE of a Z-functional is
@@
Let $\omega(X)=f_{Q_{X}}(X)/f_{P_{X}}(X)$ be the density ratio of the shared characteristic $X$ between the two populations. The weighting estimator of (11) estimates the two parts of (11) by re-weighing the source population to match the target distribution $Q_{X}$:
@@
Here, $\widehat{Q}_{O,X}^{0}$ represents samples from the source and target distributions, and $\widehat{\omega}$, $\widehat{\mu}$, and $\widehat{\psi}$ are estimates of $\omega$, $\mu$, and $\psi(Q_{O,X}^{0})$, respectively, say by parametric or semi-parametric methods. We note that the estimation of the density ratio $\omega$ is often based on balancing approaches or selection models; see Section E.7 of the Supplement for examples.
@@
For example, for the target mean $\psi^{{\textnormal{mean}}}$ with $s(O,X,\psi^{{\textnormal{mean}}})=O-\psi^{{\textnormal{mean}}}$, the weighting estimator and the regression estimator for its SLOPE are
@@
Here, $\hat{\psi}^{{\textnormal{mean}}}$ is the estimate of the target mean under conditional exchangeability and $\widehat{\sigma}^{2}(x)$ is an estimate of $\sigma^{2}(x)$. A classic way is to estimate $\widehat{\mu}(X)$ with least squares and $\widehat{\sigma}^{2}(X)$ with a log-variance linear model from weighted least squares (Harvey,, 1976; Carroll and Ruppert,, 1982; Davidian and Carroll,, 1987).
4.3 “Sensitivity” of Our Results to Sensitivity Model (3)

### v2

@@ 3.2 Example: SLOPE for the Mean and Robust Study Designs
Then SLOPE for the mean is the derivative of $\psi^{{\textnormal{mean}}}(Q_{O,X}^{\gamma})$ with respect to $\gamma$, evaluated at $\gamma=0$. In principle, researchers can compute this derivative using single-variable calculus and Theorem 1 states the regularity conditions to ensure the existence of this derivative.
Theorem 1 (SLOPE of Mean).
@@
Suppose Condition 1 holds with $s(O,\psi^{{\textnormal{mean}}})=O-\psi^{{\textnormal{mean}}}$. Then the SLOPE of the mean $\psi^{\rm mean}$ from Example 1 is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})=\textnormal{E}_{Q_{X}}\{\sigma^{2}(X)\},\text{ where }\sigma^{2}(X)=\textnormal{Var}_{P_{O\mid X}}(O\mid X).$ (6)
In words, the SLOPE of the mean is the average variability of $O$ after adjusting for $X$ in the source population and the average is taken over the target’s $Q_{X}$. When this variation is homoskedastic/constant across $X$ (i.e., $\sigma^{2}(X)=\sigma^{2}$), the mean’s SLOPE simplifies to $\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{mean}}})=\sigma^{2}$. In this case, SLOPE is only determined by the source distribution, specifically $P_{O\mid X}$; the target distribution $Q_{X}$ does not change SLOPE.
Two immediate implications follow from Theorem 1 about robust study designs. First, suppose the shared covariate $X$ explains almost all the variation in $O$ in the source population, then $\sigma^{2}(X)$ will be close to zero. Thus, SLOPE will be close to zero, meaning that the target mean will not change dramatically even if conditional exchangeability is slightly violated.
As a concrete example, consider Example 4 where the goal is to generalize the ATE by letting O=Y(1)Y(0)O=Y(1)-Y(0). If the randomized experiment in the source population suggests that the individual treatment effect is nearly constant (i.e., $Y(1)-Y(0)\approx c$ for some constant $c$), then $\sigma^{2}(X)$ will be close to zero and according to SLOPE, the ATE will not be sensitive even if conditional exchangeability is violated. In short, a near-constant treatment effect is robust in generalization. We briefly remark that Tipton and Olsen, (2018) made a similar observation in the context of generalizing ATEs in empirical works where constant treatment effects generalize better than heterogeneous ones. The main difference between Tipton and Olsen, (2018) and our work is we provide a more theoretical foundation for why constant treatment effects, or more broadly any effects where the conditional variance of O=Y(1)Y(0)O=Y(1)-Y(0) given $X$ is small, generalizes better.
Second, suppose the variation in $O$ is not well-explained by $X$ in some regions of $X$ from the source population. Then, the mean’s SLOPE can be made small by choosing a target distribution $Q_{X}$ that concentrates around a region of $X$ that shows the smallest variation in $O$ in the source population. Figure 2 provides a visual illustration of these examples and Corollary 1 formalizes these observations.
Figure 2: Illustration of designing robust studies for generalization with $\psi^{{\textnormal{mean}}}$. The x-axis represents $X$ and the y-axis represents $O$. For a formal result behind the illustrations, see Corollary 1.
Corollary 1 (Robust Study Design for Learning $\psi^{\rm mean}$).
@@
(i) Consider two source populations $P$ and $\widetilde{P}$ that satisfy $\sigma^{2}(X)\leq\widetilde{\sigma}^{2}(X)$ almost surely for a target distribution $Q_{X}$, where $\sigma^{2}(X)=\textnormal{Var}_{P_{O\mid X}}(O\mid X)$ and $\widetilde{\sigma}^{2}(X)=\textnormal{Var}_{\widetilde{P}_{O\mid X}}(O\mid X)$. Then
@@
(ii) Next, consider two target distributions $Q_{X}$ and $\widetilde{Q}_{X}$ over a common support $\mathcal{S}_{X}$ such that there exists a subset $\mathcal{S}_{X,1}\in\mathcal{S}_{X}$ that satisfies $Q_{X}(\mathcal{S}_{X,1})\leq\widetilde{Q}_{X}(\mathcal{S}_{X,1})$. If there exists a constant $c$ such that $\sigma^{2}(X)<c$ for $X\in\mathcal{S}_{X,1}$ and $\sigma^{2}(x)>c$ for $x\not\in\mathcal{S}_{X,1}$, then
@@ 3.3 Example: SLOPE for Median
Similar to the SLOPE of the mean, the SLOPE of the median also depends on the dispersion of the underlying distribution $Q_{O,X}^{0}$. Theorem 2 provides a general formula for the SLOPE of the median along with two special cases.
Theorem 2 (SLOPE of Median).
Suppose Condition 1 holds with $s(O,X,\psi^{{\textnormal{med}}})=\mathds{1}(O\leq\psi^{{\textnormal{med}}})-\mathds{1}(O>\psi^{{\textnormal{med}}})$ and $\psi^{\rm med}$ is unique, where $\mathds{1}(\cdot)$ is the indicator function which is one if the event holds and zero otherwise. Then, the SLOPE of the $\psi^{\rm med}$ from Example 2 is
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})=$ $\dfrac{\textnormal{E}_{Q_{X}}\left[F_{P_{O\mid X}}(m_{1/2}\mid X)\mu(X)\right]-\textnormal{E}_{Q_{O,X}^{0}}\left[O\mathds{1}(O\leq m_{1/2})\right]}{f_{Q_{O}^{0}}(m_{1/2})},$ (7)
where $m_{1/2}=F_{Q_{O}^{0}}^{-1}(1/2)$ is the median of $O$ on the target population under conditional exchangeability and we recall $\mu(X)=\textnormal{E}_{P_{O\mid X}}(O\mid X)$ is the conditional expectation of $O$ given $X$ in the source population.
(i) If $P_{O\mid X}$ is symmetric with respect to O and $\mu(X)=m_{1/2}$ almost surely $Q_{X}$, then (7) simplifies to
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{\textnormal{med}})=\dfrac{m_{1/2}-\textnormal{E}_{Q_{O}^{0}}\left(O\mid O\leq m_{1/2}\right)}{2f_{Q_{O}^{0}}(m_{1/2})}.$ (8)
@@
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi^{{\textnormal{med}}})=\textnormal{E}_{Q_{X}}\left[\sigma^{2}(X)\cdot\dfrac{f_{P_{O\mid X}}(m_{1/2}\mid X)}{\textnormal{E}_{Q_{X}}\left\{f_{P_{O\mid X}}(m_{1/2}\mid X)\right\}}\right].$
Compared with the SLOPE of the mean, the SLOPE of the median depends on the spread of $Q_{O,X}^{0}$ in a way that is more complicated than $\sigma^{2}(X)$. For example, in part (i) which induces symmetry, the SLOPE of the median in (8) depends on two quantities: (a) the difference between the mean and the truncated mean that is lower than the median $\left\{m_{1/2}-\textnormal{E}_{Q_{O}^{0}}\left(O\mid O\leq m_{1/2}\right)\right\}$, and (b) the inverse of the marginal density at the median $1/f_{Q_{O}^{0}}(m_{1/2})$. In essence, (a) measures the spread of $Q_{O}^{0}$ and (b) is called Tukey’s sparsity (Tukey,, 1965), which is designed to measure the inverse of the concentration of $Q_{O}^{0}$. In part (ii) when $P_{O\mid X}$ is Gaussian, the SLOPE of the median becomes a weighted average of the conditional variance $\sigma^{2}(X)$ where the weight is determined by the conditional density $f_{O\mid X}$ evaluated at the median $m_{1/2}$. This form of the median’s SLOPE resembles the mean’s SLOPE (6), which is an (unweighted) average of $\sigma^{2}(X)$. In general, whether the researcher should study the median or the mean as a measure of centrality of $O$ in their study will depend on the conditional variance $\sigma^{2}(X)$ and the shape of the tail of $P_{O\mid X}$.
@@ 4.1 Relationship Between SLOPE and IF
Another local, derivative-based measure of robustness that precedes and complements our work is Hampel, (1974)’s celebrated influence function (IF) [^1] Hampel, (1974) originally called IF the “influence curve.” In later works, others, including Hampel, referred to the influence curve as the IF due to its generalization to higher dimensions (Hampel et al.,, 2011).. In this section, we show how SLOPE is related to IF. Briefly, the IF of $\psi(\cdot)$ under $Q_{O,X}^{0}$ is a derivative-based local measure of robustness that quantifies the effect of an infinitesimal contamination at the point $(o,x)$ on the estimand $\psi(Q_{O,X}^{0})$,
@@
The term $\delta_{o,x}$ represents the dirac delta function at $(o,x)$. From the definitions, both SLOPE and IFs are local measures of robustness. The main difference is that IFs measure the local change of the target estimand due to contamination of a single data point whereas SLOPE measures the local change of the target estimand due to violation of conditional exchangeability from the entire distribution.
In addition to their definitions, we can also compare the IF and SLOPE from a geometrical perspective. Specifically, both are directional derivatives at the “origin” (i.e., $t=0$ or $\gamma=0$), but they differ in their direction. Theorem 3 shows that SLOPE is equal to IF if the IF it “tilted” towards a particular direction.
@@
Suppose either (i) $\psi$ is a Z-estimand defined in (2) and Condition 1 hold, or (ii) Conditions 2-3 in the Supplement hold. Then SLOPE can be written as
$\textnormal{SLOPE}(Q_{O,X}^{0},\psi)$ $=\textnormal{E}_{Q_{X}}\left(\textnormal{E}_{P_{O\mid X}}\left[\textnormal{IF}(O,X,\psi(Q_{O,X}^{0}))\left\{O-\mu(X)\right\}\mid X\right]\right).$ (10)
Similar to Condition 1, Conditions 2-3 are regularity conditions that ensure the chain rule under Hadamard differentiability holds. In words, (10) states that SLOPE is the expectation (over $Q_{X}$) of the conditional covariance of the IF and $O-\mu(X)$ (over $P_{O\mid X}$). As discussed in Remark 4, the term $O-\mu(X)$ is the “residual variation” of violating conditional exchangeability under the sensitivity model (3) that is unexplained by $X$. If the IF is nearly orthogonal to the residual subspace $O-\mu(X)$ based on violating conditional exchangeability, then SLOPE will be close to zero. For a Z-estimand in (2), its SLOPE will be smaller if the score function $s$ is chosen to be nearly orthogonal to the subspace spanned by $O-\mu(X)$.
In addition to the geometric interpretation of SLOPE, Theorem 3 provides a general formula to derive the SLOPE given an IF. Sections 3.2 and 3.3 derived SLOPEs for the mean and the median using their respective IFs. Also, Section C of the Supplement derives SLOPEs for scale parameters, ordinary least squares (OLS) coefficients, Pearson correlation, L-estimands and other Z-estimands using Theorem 3.
@@ 4.2 Estimation
We briefly discuss two estimators of SLOPE, the weighting estimator and the regression estimator, for Z-estimands in Example 3. In short, estimation of SLOPE follows from existing theory on M-estimation (e.g., Chapter 5 of Van der Vaart, (2000)). Details behind the estimators, including implementation, asymptotic properties, and simulations to assess their finite-sample performance, are in Sections E - G of the Appendix. Our data analysis in Section 5 uses the regression estimator below due to better finite-sample performance.
Suppose we have $n_{q}$ independent and identically distributed (i.i.d.) samples $X_{i}\sim Q_{X}$ and another independent $n_{p}$ i.i.d. samples $(X_{i},O_{i})\sim P_{O,X}$ with $i=1,\ldots,n_{q},n_{q}+1,\cdots,n_{p}+n_{q}=n$. Under Condition 1, the SLOPE of a Z-estimand is
@@
Let $\omega(X)=f_{Q_{X}}(X)/f_{P_{X}}(X)$ be the density ratio of $X$ between the two populations. The weighting estimator of (11) re-weighs the source samples to match the target distribution $Q_{X}$:
@@
The term $\widehat{Q}_{O,X}^{0}$ represents samples from the source and target distributions, and $\widehat{\omega}$, $\widehat{\mu}$, and $\widehat{\psi}$ are estimates of $\omega$, $\mu$, and $\psi(Q_{O,X}^{0})$, respectively, say by parametric or semi-parametric methods. For example, $\widehat{\mu}$ can be from OLS, $\widehat{\omega}$ can be estimated from balancing methods or selection models (see Section E.7 of the Appendix), and $\widehat{\psi}$ can be estimated by a Z-estimator with data from $\widehat{Q}_{O,X}^{0}$.
@@
As a concrete example, consider the target mean $\psi^{{\textnormal{mean}}}$ with $s(O,X,\psi^{{\textnormal{mean}}})=O-\psi^{{\textnormal{mean}}}$. The weighting estimator and the regression estimator of the mean’s SLOPE are
@@
Here, $\hat{\psi}^{{\textnormal{mean}}}$ is the estimate of the target mean under conditional exchangeability and $\widehat{\sigma}^{2}(x)$ is an estimate of $\sigma^{2}(x)$. A simple way is to estimate $\widehat{\mu}(X)$ is via OLS and $\widehat{\sigma}^{2}(X)$ is with a log-variance linear model from weighted least squares (e.g., Harvey, (1976); Carroll and Ruppert, (1982); Davidian and Carroll, (1987)).
4.3 Extending SLOPE to Other Types of Sensitivity Analysis